OpenStudy (anonymous):

Physics Question!: Can you help me find the velocities in the x and y direction when a ball is thrown a distance of 9.14 meters and the amount of time it took to land in another person's hand is 1.37 seconds. Use kinematics please.

OpenStudy (anonymous):

vel. at which time. tell me the way of throwing. is it horizontally or vertically or parabolic.

OpenStudy (anonymous):

initial velocities, and parabolic.

OpenStudy (anonymous):

here the horizontal distance(R) it travel is =initial horizontal vel.(Ux) * total time taken(T) R =Ux*T Ux=R/T=9.14/1.37 =6.67 m/s now the total vertical displacement is zero as it start from the ground & end with it. vertical displacement(H) =Uy*T-1/2(a*T^2) (Uy=initial vertical vel. ,a=vert. acc. = g=9.8) H=Uy*T -1/2( g*T^2) 0 =Uy*1.37 -0.5*9.8*1.37*1.37 Uy =9.19/1.37= 6.713 m/s so Ux=6.67 m/s Uy= 6.713 m/s